Mason Company has two manufacturing departments—Machining and Assembly. The company considers all of its manufacturing overhead costs to be fixed costs. It provided the following estimates at the beginning of the year as well as the following information with respect to Jobs A and B:

Estimated Data Machining Assembly Total
Manufacturing overhead $ 2,250,000 $ 225,000 $ 2,475,000
Direct labor-hours 15,000 150,000 165,000
Machine-hours 150,000 10,000 160,000
Job A Machining Assembly Total
Direct labor-hours 5 10 15
Machine-hours 11 2 13
Job B Machining Assembly Total
Direct labor-hours 4 5 9
Machine-hours 12 3 15

Required:

1. If Mason Company uses a plantwide predetermined overhead rate with direct labor-hours as the allocation base, how much manufacturing overhead cost would be applied to Job A? Job B?
2. Assume that Mason Company uses departmental predetermined overhead rates. The Machining Department is allocated based on machine-hours and the Assembly Department is allocated based on direct labor-hours. How much manufacturing overhead cost would be applied to Job A? Job B? (Round your intermediate calculations and final answers to 2 decimal places.)
Answer #1

1) Overhead rate = 2475000/165000 = 15 per DLH

Job A Job B
Manufacturing overhead applied 15*15 = 225 9*15 = 135

2) Machining Department rate = 2250000/150000 = 15 per MH

Assembly Department rate = 225000/150000 = 1.5 per DLH

Job A Job B
Manufacturing overhead applied (11*15+10*1.5) = 180 (12*15+5*1.5) = 187.50